a wooden cube floats in water partially immersed,when 200g weight is put on it. it is further immersed by 2cm.the lenght of the side of the cube is?

Grade:10

1 Answers

Arun
6 years ago
Dear student
 
Let length of one edge = x
Vol. of cube = L*B*W
Vol. of cube = x*x*x
Vol. of cube = x^3
Vol. of cube = Area * height.
Vol. of cube = x² * 2
Vol. of cube = 2x² cm^3
200 mg = 2x²
x² = 200 mg / 2
x = | ± √100 |
x = 10 cm
 
 
Regards
Arun (askIITians forum expert)